OpenStudy (anonymous):

Gold in the form of solid cylinders will be molded into solid blocks. Each cylinder has a diameter of 6.19 cm and a height of 6 cm. Each solid block will have a volume of 360 cm3. How many cylinders of gold will it take to mold one block of gold?

OpenStudy (anonymous):

well area of a cylinder is A=(pi*r^2)h so you would have to find the area of the cylinder and divide 360 by it go to the last whole number like if its something like 6.8 it can only make 6 full cylinders

OpenStudy (anonymous):

so when we substitute A=(3.14*6.19^2)(6)

OpenStudy (anonymous):

sorry not 6.19, 3.095 due to the fact we need the radius not the diameter

OpenStudy (anonymous):

A=about 180.5 and 360/180.5=1.99 so it would take 2 cylinders to make a block of gold

OpenStudy (anonymous):

i just reread the question so in this case you would round up in order to get enought area from the cylinders to make a gold block not the other way around